On June 30 2018 Georgia Atlantic inc leased warehouse equipment from builders inc. the lease agreement calls for Georgia Atlantic to make semiannual lease payments of $562907 over a three year lease term, payable each June 30 and December 31 with the first payment at June 30 2018. Georgia Atlantic’s incremental borrowing rate is 10%, the same rate builders used to calculate lease payment amounts. Builders used to calculate lease payment amounts. Builders manufactured at a cost of $2.5 million.

REQUIRED:

1. Determine the price at which builders is selling the equipment (present value of the lease payment) at June 30 2018 ( to the nearest $000)

2. What amounts related to the lease would builders report in its balance sheet at December 31 2018( ignore)

3. What amounts related to the lease would builders report in its income statement for the ended December 31, 2018( ignore taxes)
Answer #1

Solution 1:

Semiannual lease payment to be received = $562907

Total semiannual payments to be received = 3*2 = 6

Incremental borrowing rate = 10%, 5% semiannual

Price at which Builders is “selling” the equipment = Semi Annual lease payments * Cumulative PV Factor of annuity due for 8 periods at 6%

= $562907 * 5.32948 = $3,000,000

Solution 2:

Semiannual payment received on 30.06.2018 = $562907

Pretax amount of liability on 30.06.2018 = ($3,000,000 - $562907) = $2,437,093

Interest revenue for 31.12.2018 = $2,437,093 * 5% = $121,855

Semiannual lease payment received on 31.12.2018 = $562907

Pre tax amount for receivables on December 31, 2018 = $2,437,093 + $121,855 - $562907 = $1,996,041

Solution 3:

Pre tax amount of interest Revenue Builder Inc. reports in its income statement = $2,437,093 * 5% = $121,855

Pre tax amount of revenue Builder Inc. reports in its income statement = $3,000,000

Pre tax amount of cost of goods sold to be reported by Builder Inc. in its income statement = $2,500,000

Builders report in its income statement for the ended December 31, 2018 = Revenue + Interest - Cost of goods sold

= $3,000,000 + $121,855 - $2,500,000 = $621,855